OpenStudy (anonymous):

Use dimensional analysis to determine the number of kilograms in 62 pounds. (Hint: 1 kg = 2.2 lbs) 136.4 kg 0.04 kg 281.8 kg 28.2 kg

OpenStudy (anonymous):

Usually "dimensional analysis" is a fancy word for "rounding things to nice numbers and gussing."

OpenStudy (anonymous):

Like, 2.2 * 62 is ugly. 2 * 60 is easy. What number there is closest to 120?

OpenStudy (anonymous):

Er, wait, divide, don't multiply. 60/2. But same idea!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Directrix wht is your opinion on this

Directrix (directrix):

62/2.2 = 28.123 kg Yes? @Starr_Dynasty

OpenStudy (anonymous):

TO WIKIPEDIA WE GO http://en.wikipedia.org/wiki/Dimensional_analysis

OpenStudy (anonymous):

That article does not help at all.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i believe that the answer is A its pretty much logical to me

OpenStudy (anonymous):

You multiplied. You have to divide.

OpenStudy (anonymous):

60/2 = 30 Which one is closest to 30?

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

Use dimensional analysis to determine the number of kilograms in 62 pounds. (Hint: 1 kg = 2.2 lbs) 136.4 kg 0.04 kg 281.8 kg 28.2 kg

OpenStudy (anonymous):

You don't think 28.2 is close to 30?

OpenStudy (anonymous):

wait D

OpenStudy (anonymous):

Ok. :P

OpenStudy (anonymous):

sorry i was seeing more than one digits

OpenStudy (anonymous):

All "dimensional analysis" really means, at least as I've ever heard it used, is to reduce everything to quantities you can grasp with simple mental math and make sure something is logical.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

So, like that, don't try to work out 62/2.2, do 60/2 in your head.

Directrix (directrix):

That's what I said: 62/2.2 = 28.123 kg Yes?

OpenStudy (anonymous):

ok i understand

OpenStudy (anonymous):

u didnt explain of how u got your answer it just looked like numbers pretty much

OpenStudy (anonymous):

Actually doing the division doesn't count! :P